Suppose that the supply of barley is given by the supply function : Qs= 500+5p+2r

where Qs is the quantity of barley supplied, P is the price of barley, and r is total rainfall in inches.

Suppose that the demand function for barley is given by: Qd= 700-8p

d) now assume that consumer preferences have changed, and the new demand curve for barley is Qd= 700-6p

after a large increase in rainfall, will the equilibrium quantity be higher or lower with the new demand function as opposed to the old demand function why? use a graph to help your explanation  
Answer #1

Note: Change in demand for Q = 700-8P to Q = 700-6P is an increase in demand and shift demand curve go to right.